OpenStudy (anonymous):

integral of sin (3x) cos^6 (3x)

OpenStudy (aum):

Let u = cos(3x)

OpenStudy (aum):

\[ \int \sin(3x)\cos^{6}(3x)dx \\ \text{Let } u = \cos(3x) \\ du = -3\sin(3x)dx \\ \sin(3x)dx = -\frac 13du \\ \int \sin(3x)\cos^{6}(3x)dx = \int -\frac 13 u^6du = -\frac 13 \frac{u^7}{7} = \\ -\frac{1}{21}\cos^7(3x) + C\\ \]

OpenStudy (anonymous):

Thanks :)

OpenStudy (aum):

you are welcome.